How do you write m^6-16m^3+64=0 in quadratic form?
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


Quadratic form means the highest power of the variable is 2.


In problems like this one note that m^6 = (m^3)^2. So make some other variable, say u, and set u=m^3. Then this equation becomes:


u^2-16u+64 = 0